# Geonhwa Jeong

### v2

@@ 5.3 Analysis of TASD
The left plot of Figure 16 shows the impact of network-wise TASD-W on the top-1 accuracy of unstructured sparse ResNet50 (95% sparsity). We applied network-wise TASD-W with N:4, N:8, and N:16 structured sparsities. For example, the network-wise TASD-W with 2:4 uses one TASD term with the 2:4 pattern to the weights of all convolution and fully-connected layers in the sparse ResNet50. Since TASD is a lossy method as shown in Section 3.2, aggressive TASD series approximation can result in a notable accuracy drop. Among different N:4, N:8, N:16 options, we found that 3:4, 5:8, 10:16 is the most aggressive approximation among the available options while meeting the 99% accuracy requirement. Especially, using network-wise TASD-W 5:8 (4:8 + 1:8 for TTC-VEGETA) and power gating for sparse activations, compared to the dense baseline, we observe it achieves 24% and 53% reduction in cycle and energy respectively, thus reducing 75% EDP for Sparse ResNet50.
Using different TASD series configurations for different layers is more effective as it can adjust the aggressiveness for each layer. To choose a TASD configuration per layer, we use the sparsity-based selection method that we introduce in Section 4.2. By changing the hyperparameter alpha, we are able to adjust the aggressiveness of our approximation method. As layer-wise TASD-W can be adaptive to each layer, the overall approximation can be applied more aggressively. As a result, compared to the dense baseline, we observe 47% and 61% reduction in energy and cycle respectively, thus reducing 83% EDP for Sparse ResNet50.
In the right plot of Figure 16, we show the top-1 accuracy when network-wise and layer-wise TASD-A is applied with different TASD series. Similar to TASD-W, layer-wise TASD-A is more effective than network-wise TASD-A. However, the accuracy loss due to approximation shows up with a much smaller approximated sparsity. As shown earlier in  Figure 8, the sparsity degree is larger in weights compared to that in activations for sparse ResNet50. Thus, the same TASD series drops a larger portion of non-zeros in TASD-A than TASD-W, incurring a higher loss of accuracy.
TASD on more DNN models.
To further investigate the impact of TASD-W on different sparse DNNs, we applied layer-wise TASD-W on different Sparse ResNet and VGG families with a requirement to maintain 99% of the original accuracy. We use the pre-trained unstructured sparse models from SparseZoo [45]. Across different sparse ResNet models and VGG models, TASD-W reduced 49% MAC operations while maintaining 99% accuracy, as shown in Figure 17.
Figure 17: Layer-wise TASD on more DNN models. Left: TASD-W. Right: TASD-A.
On the other hand, to understand the potential impact of TASD-A on other DNN models, we applied TASD-A on various models including both convolution networks and a transformer-based network, as shown in Figure 17. We use the pre-trained dense models from TorchVision [1] and Huggingface for this evaluation and we use the requirement to meet 99% of the original accuracy. We observe that the layerwise TASD-A is effective for various models and achieves 32% reduction in MACs for other models on average.
5.4 Comparison against structured sparse accelerators
5.5 Energy overhead due to TASD
@@
Figure 19 shows the energy breakdown for a representative layer from sparse ResNet50 for dense TC and TTC-VEGETA with a TASD configuration of 4:8+1:8. TTC-VEGETA exploits sparsity and saves energy consumption at all levels of the architecture, which saves 55% energy over the dense TC. Moreover, the decomposition-aware dataflow in Section 4 minimizes decomposition overheads by accessing the RF (with C reuse) and SMEM (with B reuse) instead of accessing DRAM.
5.6 Area overhead of the TASD unit
We also measured the area overhead to support TASD on top of the existing structured sparse HW, i.e. the TASD unit, through RTL prototyping and synthesis with Nangate 15nm Technology Library. We observe up to 2% of the area for all PEs as TASD units are composed of simple comparator trees.
@@ 6.1 SW techniques for structured sparse HW
Since the structured sparse (2:4) accelerator is available in off-the-shelf NVIDIA GPUs with Ampere [46] and Hopper [48] architecture, ample work has focused on exploiting structured sparse acceleration for DNNs.
Solutions with fine-tuning. DominoSearch [62] proposed a method to find layer-wise N:M sparsity during training. Optimal N:M [7] enforces the structured pattern during training, but with a minimum-variance-based pruning method, and applies the structured pattern to input activations. This line of work is orthogonal to our work as we focus on approximating unstructured sparsity without fine-tuning. Fine-tuning will increase the benefit of TASD, as more aggressive approximation can now maintain the same model accuracy. Doping [63] uses an extremely sparse matrix in addition to a compressed matrix derived from Kronecker products to improve the quality of the model, but unlike TASD, it uses the extra extremely sparse matrix to give additional freedom during the training, not for the approximating the given sparse matrix.
Solutions without fine-tuning. SparseTIR [68] introduces composable formats and transformations for sparse compilation of deep learning workloads. However, they have not considered approximating sparse tensors and accelerating DNNs using structured sparse hardware. Another work [52] shows permuting channels in the weight tensors can recover accuracy easily when training N:M sparse networks. TASD is compatible with channel permutation, and we believe combining these two orthogonal techniques will further improve the accuracy of decomposed models with aggressive approximation.
@@ 6.2 HW support for sparse DNNs
Sparsity support for DNN inference. Different architectures have been proposed to support for weight sparsity [6, 28, 46, 17], for activation sparsity [26], and more recently, for both [64, 23, 65]. As mentioned earlier, these accelerators can be broadly classified as structured sparse or unstructured sparse HW. Unstructured sparse HW provides native support for any sparsity pattern but is more costly to build; structured sparse HW is efficient but requires model fine-tuning. TASD bridges the gap in this area by providing an unstructured sparse interface while only requiring structured sparse HW.
Sparsity support for DNN training. Since weight tensors are mostly dense during training, prior work has focused on activation and gradient sparsity during DNN training. The simpler support is to compress sparse activation and gradient, such as CompressDMA [57] and ZComp [3]. These techniques save data movements and memory requirements, but not overall compute. The more complex techniques target reducing computation during training, such as TensorDash [39] and SAVE [18]. However, they need to give up data movement savings for better support for sparse tensor transposition during training. TASD can be used to approximate sparse activations and gradients, too. As prior work [33, 57] has shown that the degree of activation and gradient sparsity stays stable, TASD-A can approximate these dynamic tensors during runtime to enable the efficient structured sparse HW support for sparse training. We leave this to future work.
@@ 7 Conclusion
Sparse DNN model developers prefer to induce unstructured sparsity for expressibility, while sparse DNN hardware designers prefer to support structured sparsity for HW efficiency. This mismatch of the desired sparsity pattern prevents sparse DNN acceleration from being widely adopted in practice.
To close the gap, we introduce TASD, a method that approximates an unstructured sparse tensor with a series of structured sparse tensors. Next, we propose a framework, TASDER, which finds TASD configuration for each DNN layer to accelerate off-the-shelf sparse and dense DNNs. To maximize the benefit of TASD, we propose a simple architectural extension and dataflow on top of structured sparse accelerators. TASD improves EDP by up to 83% and 74% on average, while maintaining 99% of the model accuracy without any fine-tuning.

### v3

@@ 5.3 Analysis of TASD
The upper plot of Figure 14 shows the impact of network-wise TASD-W on the top-1 accuracy of unstructured sparse ResNet50 (95% sparsity). We applied network-wise TASD-W with N:4, N:8, and N:16 structured sparsities. For example, the network-wise TASD-W with 2:4 uses one TASD term with the 2:4 pattern to the weights of all convolution and fully-connected layers in the sparse ResNet50. Since TASD is a lossy method, aggressive TASD series approximation can result in a notable accuracy drop. Among different N:4, N:8, N:16 options, we found that 3:4, 5:8, 10:16 is the most aggressive approximation among the available options while meeting the 99% accuracy requirement. Especially, using network-wise TASD-W 5:8 (4:8 + 1:8 for TTC-VEGETA) and gating the compute units for sparse activations, compared to the dense baseline, we observe it achieves 24% and 53% reduction in cycle and energy respectively, thus reducing 75% EDP for Sparse ResNet50.
Using different TASD series configurations for different layers is more effective as it can adjust the aggressiveness for each layer. To choose a TASD configuration per layer, we use the sparsity-based selection method that we introduce in section 4. By changing the hyperparameter, alpha, we are able to adjust the aggressiveness of our approximation method. As layer-wise TASD-W can be adaptive to each layer, the overall approximation can be applied more aggressively. As a result, compared to the dense baseline, we observe 47% and 61% reduction in energy and cycle, respectively, thus reducing 83% EDP for Sparse ResNet50 as shown in Figure 12.
In the lower plot of Figure 14, we show the top-1 accuracy when network-wise and layer-wise TASD-A is applied with different TASD series. Similar to TASD-W, layer-wise TASD-A is more effective than network-wise TASD-A. However, the accuracy loss due to approximation shows up with a much smaller approximated sparsity. As shown earlier in  Figure 6, the sparsity degree is larger in weights compared to that in activations for sparse ResNet50. Thus, the same TASD series drops a larger portion of non-zeros in TASD-A than TASD-W, incurring a higher loss of accuracy.
5.4 Energy and area overhead due to TASD
@@
Figure 16: TASD-W on NVIDIA RTX3080 GPU with Sparse ResNet34.
Figure 15 shows the energy breakdown for a representative layer from sparse ResNet50 for dense TC and TTC-VEGETA with a TASD configuration of 4:8+1:8. TTC-VEGETA exploits sparsity and saves energy at all levels of the architecture, which saves 55% energy over the dense TC. Moreover, the decomposition-aware dataflow in section 4 minimizes decomposition overheads by accessing the RF (with C reuse) and SMEM (with B reuse) instead of accessing DRAM.
We measured the area overhead to support TASD on top of the existing structured sparse HW (the TASD units) through RTL prototyping and synthesis with Nangate 15nm Technology Library. We observe up to 2% of the area for all PEs as TASD units are composed of simple comparator trees.
5.5 Using TASD on a real system
To validate the effectiveness of TASD, we evaluate TASD on a real system. First, we use our framework, TASDER, to find the TASD-W configuration that meets the accuracy constraint with an unstructured sparse ResNet34 model from SparseZoo Neuralmagic (2023). For the accuracy evaluation, we use ImageNet as the dataset. Next, we export the model using ONNX format. To properly measure the inference latency of DNNs exported with ONNX format, we use TensorRT, the state-of-the-art deep learning inference runtime maintained by NVIDIA. We build TensorRT engines for both the baseline model and the model with TASD-W exported in the previous step. Finally, we execute the engine with (3, 224, 224) inputs, with a batch size of 32-128, and measure the latency for each model. We achieved 28%/39% speed-up compared to the dense model execution with 0.9%/1.5% accuracy drop as shown in Figure 16. We also evaluated TASD-W with ResNet50, ResNet101, BERT, and observed up to 25% speed-up.
@@ 6.1 SW techniques for structured sparse HW
Solutions with fine-tuning. DominoSearch Sun et al. (2021) proposed a method to find layer-wise N:M sparsity during training. Optimal N:M Chmiel et al. (2022) enforces the structured pattern during training and applies the structured pattern to input activations. This line of work is orthogonal to our work as we focus on approximating unstructured sparsity without fine-tuning. Fine-tuning will increase the benefit of TASD, as more aggressive approximation can now maintain the same model accuracy. Doping Thakker et al. (2021) uses an extremely sparse matrix in addition to a compressed matrix derived from Kronecker products to improve the quality of the model, but unlike TASD, it uses the extra extremely sparse matrix to give additional freedom during the training.
Solutions without fine-tuning. SparseTIR Ye et al. (2023) introduces composable formats and transformations for sparse compilation of DL workloads. However, they have not considered approximating sparse tensors and accelerating DNNs using structured sparse hardware. Another work Pool & Yu (2021) shows permuting channels in the weight tensors can recover accuracy easily when training N:M sparse networks. TASD is compatible with channel permutation, and we believe combining these two orthogonal techniques will further improve the accuracy of decomposed models with approximation.
@@ 6.2 HW support for sparse DNNs
Sparsity support for DNN inference. Different architectures have been proposed to support for weight sparsity Chen et al. (2019); NVIDIA (2020a); Jeong et al. (2021); Gondimalla et al. (2023), for activation sparsity Jang et al. (2021), and more recently, for both Wang et al. (2021); Huang et al. (2023); Wu et al. (2023). As mentioned earlier, unstructured sparse HW provides native support for any sparsity pattern but is more costly to build; structured sparse HW is efficient but requires model fine-tuning. TASD bridges the gap by providing an unstructured sparse interface while only requiring structured sparse HW.
Sparsity support for DNN training. Since weight tensors are mostly dense during training, prior work has focused on activation and gradient sparsity during DNN training. A simpler support is to compress sparse activation and gradient, such as CompressDMA Rhu et al. (2018) and ZComp Akin et al. (2019). These techniques save data movements and memory requirements, but not overall compute. The more complex techniques target reducing computation during training, such as TensorDash Mahmoud et al. (2020) and SAVE Gong et al. (2020). However, they need to give up data movement savings for better support for sparse tensor transposition during training. TASD can potentially be used to approximate sparse activations and gradients, but we leave this to future work.
@@ 7 Conclusion
Sparse DNN model developers prefer to induce unstructured sparsity for expressibility, while sparse DNN hardware designers prefer to support structured sparsity for HW efficiency. This mismatch of the desired sparsity pattern between different parties prevents sparse DNN acceleration from being widely adopted in practice. To close the gap, we introduce TASD, a method that approximates an unstructured sparse tensor with a series of structured sparse tensors. Next, we propose a framework, TASDER, which finds TASD configuration for each DNN layer to accelerate sparse and dense DNNs. To maximize the benefit of TASD, we propose a simple architectural extension and dataflow on top of structured sparse accelerators. TASD improves EDP up to 83% and 74% on average, while maintaining 99% of the model accuracy without any fine-tuning. We also show it achieves up to 39% performance improvement on a real system evaluation.
Acknowledgements
We thank Jeff Pool, Angshuman Parashar, Michael Pellauer, Qijing Huang, Joel Emer, Christopher Hughes, Hyesoon Kim, and Vivek Sarkar for their valuable feedback and insightful discussions.
Appendix A Additional details of TASD
A.1 An analysis of TASD with synthetic data
We generate a synthetic matrix with dimensions of 128$\times$128 and densities ranging from 0.1 to 0.75. We explore three TASD series in this experiment; 1) using one term with 2:4, 2) using two terms with 2:4 and 2:8, and 3) using three terms with 2:4, 2:8, and 2:16. To consider various distributions, we tested two different distributions, a uniform random distribution between 0 and 1 and a normal distribution with a mean of 0 and a standard deviation of $\frac{1}{3}$. Figure 17 shows the results with matrices generated using the normal distribution.
Error Behavior: The first trend we observe is that the error gets smaller as the approximated sparsity gets lower since it is likely to drop fewer non-zeros with a more conservative approximation. Second, for the same approximated sparsity and the block size (M), the error gets smaller as the matrix $A$ gets sparser. Given the same TASD configuration, the sparser matrices would drop fewer non-zeros using the same TASD configuration as shown in Figure 17, thus resulting in a smaller error. Third, with the same sparsity of matrix $A$ and approximated sparsity, the N:4 configuration causes a larger error than the N:8 configuration (such as 1:4 and 2:8), since the expressiveness of the N:8 pattern is higher. Finally, given any unstructured sparse tensor, we can limit the error of matrix multiplication by conservatively selecting the TASD configuration, while maximizing the compute reduction. This optimization thus becomes the key to leveraging TASD for accelerating sparse DNN models with structured sparse hardware. In terms of the number of effectual computations, using a sparser TASD configuration is favorable to approximate the matrix multiplication while it would cause a higher error if not chosen cautiously.
Appendix B Additional Experiments
In this section, we provide additional experiments to understand the impact of TASD.
B.1 Comparison to structured sparse accelerators
B.2 TASD on more DNN models
To further investigate the impact of TASD-W on different sparse DNNs, we applied layer-wise TASD-W on different Sparse ResNet and VGG families with a requirement to maintain 99% of the original accuracy. We use the pre-trained unstructured sparse models from SparseZoo Neuralmagic (2023). Across different sparse ResNet models and VGG models, TASD-W reduced 49% MAC operations while maintaining 99% accuracy, as shown in Figure 20.
Figure 20: Layer-wise TASD on more DNN models.
On the other hand, to understand the potential impact of TASD-A on other DNN models, we applied TASD-A on various models including both convolution networks and a transformer-based network, as shown in Figure 20. We use the pre-trained dense models from TorchVision PyTorch (2023) and Huggingface for this evaluation and we use the requirement to meet 99% of the original accuracy. We observe that the layerwise TASD-A is effective for various models and achieves 32% reduction in MACs for other models on average.